Consider the following statements as they apply to crystal structure. Decide if each is true or false. v Planes with Miller indices (101) are parallel to the z-axis of the crystal If gold crystallizes in a face-centered cubic lattice, each unit cell would contain four gold atoms Planes described with Miller indices (220) have M-8. ▼ There are seven crystal systems: cubic, tetragonal, orthorhombic, monoclinic, triclinic, hexagonal, and trigonal.
Answer #1

Consider the following statements as they apply to crystal structure. Decide if each is true or false. False▼ Planes with Mil

explanation :

(101) plane is parallel to y - axis but not z -axis. becuase middle y = 0

unit cell is fundamental unit not space lattice. space lattice is made of unit cell like brocks in a wall

(220) , M = sqrt (2^2+ 2^2 + 0) = sqrt (8)

not simple 8

remaining two are correcct